OpenStudy (anonymous):

A family of 4 goes to a restaurant and sits at a rectangular table with 4 chairs. How many different seating arrangements are possible for this family?

OpenStudy (ash2326):

@watermelon26 there are 4 members and 4 seats. This is a problem of permutation First member can choose from 4 chairs, so he/she can sit in 4 ways Second member will have to choose from 3 chairs so 3 ways similarly third can choose from 2=> 2 ways for the fourth member=>only 1 way so total no. of ways \[4 \times 3\times 2\times 1=4!=24\]

OpenStudy (ash2326):

@watermelon26 did you understand?